OpenStudy (anonymous):

The temperature outside changed from 78°F to 48°F over a period of six days. If the temperature changed by the same amount each day, what was the daily temperature change?

OpenStudy (lgbasallote):

(78-48)/6 = 30/6 = 5 degrees Fahrenheit per day

OpenStudy (lgbasallote):

-5 degrees..cause it's decreasing